# Towards User-Focused Research in Training Data Attribution for Human-Centered Explainable AI

### v1

@@ 4.1.1. Participants
To get a realistic impression of user needs for TDA explanations in practice, we target participants who work with ML applications outside of academia. Our inclusion criteria are: Participants should (1) have at least one month of experience in working with ML systems and (2) work in a high-risk application area according to the EU AI Act (Parliament, 2023) (e.g., health care, law enforcement, complete list in Appendix A). This criterion serves to find users who are likely to use explanations, as these areas are subject to further regulations (Guidotti et al., 2018; Doshi-Velez and Kim, 2017). Recruiting participants poses a challenge, especially in high-risk application areas. Hence, we use purposive sampling (Guest et al., 2023) and approach participants from the authors’ network. We recruit 10 participants from various domains and degrees of experience (see Table 1).
| ID | Location | Domain | Type | Job experience / with ML | Type of ML |
| P1 | DE | HR | User | 3 years / 1 months | Chatbot |
| P2 | US | AV | Developer | 2 years / 7 years | Prediction model |
| P3 | NL | TC | Developer | 3 years / 5 years | Prediction model |
| P4 | FI | CV | Developer | 4 years / 6 years | Prediction model |
| P6 | CH | Health | User | 2 years / 2 years | Prediction model |
| P7 | NL | Health | Developer | 1 year / 3 years | Prediction model |
| P8 | BE | Health | Developer | 2 years / 6 years | Prediction model |
| P9 | PK | Health | Developer | 5 years / 2 years | Prediction model |
| P10 | DE | HR | User | 3 years / 1 year | Chatbot |
| P11 | DE | Health | User | 10 years / 6 years | Clustering, Chatbot |
@@ 4.1.2. Interview process
The interviews were conducted during June – September 2023, either in person or remotely through video call using Zoom 111https://zoom.us/. All interviews are one-on-one conversations in English, except for P10 in German. Participants were first briefed about the purpose of the study and data processing. Upon receiving informed consent, we began the interview recording. In general, the interviews lasted between 30 and 60 minutes. Each interview addresses the following topics to answer RQ1 - What do users need TDA explanations for?:
Job-related information. Perspectives may vary between different domains, levels of seniority and experience with the ML tool. This information provides user context and will help in differentiating user groups.
Interviewee’s workflow with ML systems. By asking about the workflow with the ML tool, we wish to understand the patterns of usage and challenges participants encounter. We collect this information to identify challenges that could potentially addressed by TDA.
Perspectives on training data. Since we investigate TDA explanations, we explicitly ask participants about the role training data plays in their tasks. This indicates what type of data-centric information is relevant to the participant’s work and for which tasks.
Perspectives on XAI and TDA. We address the participant’s perspectives on XAI and particularly on TDA to understand current explanation usage scenarios and explore participant’s opinions about TDA, a data-centric approach to XAI.
@@ 4.1.3. Data analysis
We analyse the interview data and extract relevant information for answering RQ1. The interviews are transcribed automatically using Whisper (Radford et al., 2023) and cleaned up manually by the authors. The transcript is then pseudonymised. We translated P10’s German transcript to English using DeepL 222https://www.deepl.com/translator. We then analyse the transcripts through an inductive thematic analysis (Braun and Clarke, 2012) by two coders, where the codes are directly annotated in the transcript. The analysis is iterative: The interview transcript of P1 is first jointly analysed in an initial coding workshop that resulted in an initial set of themes. Afterwards, the coders independently code five transcripts, expanding on the themes found in the initial analysis. During an intermediate coding workshop, agreements and disagreements between the coder’s themes are discussed. The workshop resulted in a merged definition of the themes that are used for the remaining transcripts. At the intermediate coding workshop, the interrater agreement is 77.3% measured by the percentage of agreement participants coded to themes. The final coding workshop takes place after both coders reviewed the remaining transcripts. The final interrater agreement is 80.3%.
4.2. Interview findings
The figure depicts the resulting themes from the coding analysis of the interview transcripts. There are 28 themes in total, spread over six theme areas. The theme areas are larger boxes with rounded edges, encompassing the themes denoted in square boxes (intended to mimic sticky notes). Themes related to training data and TDA are marked with a star. The themes areas with themes in parentheses are: Role of ML systems (Work assistant, decision support, automation), workflow with ML systems (Collab with domain experts, standard model choices, data-focused debugging process*, data-quality checks*), Role of training data (Most important variable in a model*, interesting but no capacity*, not interesting*, increasing training data as a bug fix*, training data artifacts as main bug causes*), challenges in working with ML systems (trust calibration, resource constraints, misuse due to lacking know-how, validation and evaluation*, data quality issues*, stochastic nature of ML models*), use of feature attribution XAI in practice (XAI is not used, XAI is little used, understand real-world phenomena, get customer buy-in, sanity check for model reasoning, model development*), user perspectives on data-centric XAI (usefulness of XAI is use-case dependent*, XAI is useful*, TDA complementary to feature attribution*, different notions of interesting training samples*).
The interview study aims to provide insights into the potential usefulness of TDA, answering RQ1 - What do users need TDA explanations for? The analysis yields six theme areas, which are common across all participants (see light grey boxes in Figure 2): the role of the ML system as an application, the workflow of participants with ML systems, the role of training data for participant’s work, the main challenges in working with ML systems, the use and role of explanation tools, and perspectives and opinions about data-centric explanations like TDA. We elaborate on each area below and provide a full overview in Appendix B:
4.2.1. Role of ML systems.
Our participants work with different types of ML models across several domains and in different positions (see Table 1). We find that ML systems are used as work assistants (P1,P10,P11), decision support systems (P3,5,6,7,8,9), and for automation in systems like autonomous vehicles or machinery (P2, P4). These roles provide context for our analysis, as the needs and usage scenarios of XAI may differ based on the level of user interaction. For example, P4 develops automated computer vision models for animal population monitoring with minimal human involvement, while P10 uses a chatbot to handle smaller tasks like answering employee questions. Our analysis is set against the backdrop of these use cases.
4.2.2. Workflow with ML systems.
As ML systems take different roles, the workflows differ. We notice that ML systems are not an essential part of the workflow for system users, as opposed to developers. System users only integrate the ML system in their work when it delivers helpful suggestions and report bugs or ignore it otherwise (P6, P10). We recognise a separation of domain knowledge (collaboration with domain experts): “Because personally, I cannot know if the model is doing the correct thing […] business have to tell me” (P3). For developers whose job revolves around the ML system, we identify themes that are relevant for understanding usage scenarios of XAI and TDA. We find that the development workflow is heavily centred around data: “[What] drives your model is your data. […]” (P8) (data-focused debugging process, data quality checks). Often, standard model architectures are employed with standard hyperparameters (standard model choices) and the main work lies in data curation (P2, P4, P3, P7, P8, P9). This theme area shows that there are two distinct user groups, system users who utilise ML systems and model developers who build ML systems. They have different workflows and potential XAI use cases.
4.2.3. Role of training data.
The training data plays different roles to system users and model developers. While it is the most important variable in a model for developers (P3, P4, P8), training data is generally not interesting to system users (P1, P10). System users often do not have capacity to interact much with the ML system to verify model behaviour, even if they an interest in the underlying mechanisms exists, because they are preoccupied with their own tasks (interesting but no capacity) (P6). However, for developers, training data is a lever for model performance (increasing training data as a bug fix, training data artifacts as main bug causes) (P2, P4, P3, P7, P8, P9, P11). These insights underline that XAI needs to be intuitive, and TDA is particularly suited for model development and less for system users.
4.2.4. Challenges in working with ML systems.
In the interviews, we ask participants about challenges in their ML workflow to explore if XAI can address them. We identified six themes in this theme area. We confirm Kim et al.’s (Kim et al., 2023) observation that system users’ main challenge for working with ML systems is trust calibration (P1, P6, P8, P11). In addition, we find that system users fear they may misuse the ML system due to lacking know-how (P1, P10, P11), highlighting the need for user training to make effective use of ML systems in practice. For model developers, we identify several challenges: Data quality issues are often the root cause of model malfunction (data quality issues) (P2, P4, P3, P7, P8, P9, P11), in turn impacting model validation and evaluation. For instance, participants encounter difficulties due to missing data or absent labels (P2, P4, P3, P7, P8, P9). Additionally, resource constraints (P2, P4, P11) and the stochastic nature of ML models (P2) are pain points for the development and evaluation of ML systems. Our analysis shows that the challenges faced by system users require sociotechnical consideration, whereas challenges faced by model developers are more technical with data quality and evaluation being most pressing. We believe that the latter is a suitable usage scenario for TDA.
4.2.5. Use of XAI in practice.
@@ 4.2. Interview findings
This theme area groups six themes about the use or lack of XAI in practice. Our analysis reveals that XAI is usually not a standard part of an ML system in practice (XAI is not used, XAI is little used) (P1, P4, P6, P7, P9, P10, P11). In fact, the uses of XAI, in particular feature attribution, among our participants are limited to model development (P2, P3, P8). XAI offers explanations for per-example debugging or act as a sanity check for model reasoning (P8, P9). Unexpectedly, developer participants also use XAI to understand real-world phenomena modelled by their ML system (P3) and as a communication means to get customer buy-in for their ML systems (P8). While explanations have different purposes, we note that participants use XAI tools mainly as an out-of-the-box functionality (e.g., the SHAP library (Lundberg and Lee, 2017)). We find that implementation thresholds must be low for the adoption of XAI in practice.
4.2.6. User perspectives on TDA
@@
None of the participants were familiar with TDA, highlighting a gap between research and practice. Our analysis, based on discussions about the concept of TDA, identifies four themes reflecting participants’ views on XAI and TDA. Overall, participants find XAI useful for debugging and communication, though they noted its usefulness is use-case dependent and not always guaranteed. They expect TDA not to be an exception (P3, P6, P8, P11), but are optimistic about its potential, especially for model development. For instance, P2 mentioned that TDA could save time by identifying faulty training data. One participant suggested combining TDA with feature attribution for debugging (P3). We observed that participants have different notions of interesting training samples, with some focused on finding faulty data (P2, P11) and others on identifying samples similar to a specific one (P6, P7, P8).
In conclusion, we identify that TDA could be valuable for model debugging, leading us to focus on this user group and use case in the remainder of this work.
@@ 4.3. Defining a design space of TDA explanations
| Theme A: Data quality issues | ||||
| Theme B: Different notions of interesting training samples | ||||
| ID | Theme | Code | Description | Participants |
| #1 | A | Distribution shifts | If the difference in data distribution between training and testing data is too large, the model does not learn. This can also happen over time. | P3, P4 |
| #2 | A | Data unavailability | Missing data entries or missing labels are common problems, particularly for tabular data. | P2, P3, P7, P8, P9 |
| #3 | A & B | Wrong labels | Annotation is an expensive task subject to human error. mislabelled data is hence a possible reason for model error. | P2, P4 |
| #4 | A | Data scaling | Wrong data scaling in tabular data (e.g. using meters instead of centimetres) can lead models to learn wrong correlations. | P3 |
| #5 | A | Historical data | Outdated data that is not cleaned out of the training dataset can contaminate the training. | P9 |
| #6 | B | Outlier data points | Data outliers, i.e. mislabelled data or “unusual data points” (P3) can aid for model debugging. | P3, P8 |
| #7 | B | Decision boundary samples | Samples at decision boundary are interesting to understand the model behaviour. | P7 |
| #8 | B | High-confidence samples | Samples that the model is sure about tend to represent what the model has learnt and help in understanding the model behaviour. | P7 |
| #9 | B | Similar training samples | Training samples that are similar in features to the test sample and how the model behaves on these is helpful for understanding model behaviour. | (P6), P7, P8 |
| #10 | B | Group of samples | Inspecting multiple training samples is generally more interesting to estimate model behaviour. | P8 |
The figure depicts a three-dimensional coordinate system with categorical axes. The x-axis is named “Action defining relevance” and has the categories: remove, change label, reweight and recommend data collection. The y-axis is named “Metric” and has the categories: Loss, classification probability and ground-truth probability. The z-axis is named “Number of training samples” and has the categories: Explain f(x) with 1 training sample, and explain f(x) with N training samples.
The interview analysis reveals model developers as a key audience for TDA explanations, with model debugging as a promising use case. Before addressing RQ2 – What do users need from TDA explanations?, we translate these insights into a design space of TDA explanations in model debugging. The space shall illustrate how explanatory information about training samples and model output could look like, forming the basis for the study addressing RQ2.
Interview participants provided rich information about what they usually look out for in the training data when debugging a model. We coded this information into the following themes: Data quality issues, and different notions of interesting training samples. We present the descriptions of the corresponding codes in Table 2 and use them to identify different axes of information that could be represented by TDA explanations. Specifically, we find that different data artefacts are debugged using different actions, with users interested in various metrics and considering different numbers of training samples. From these observations, we define a three-dimensional design space (see Figure 3). In the following, we detail the reasoning behind each axis and connect it to the codes in Table 2 via the ID column.
@@ 4.3.1. Axis 1: Action defining relevance
The first axis corresponds to the actions participants propose for different types of data artefacts: Removing data, changing the label, reweighting data and collecting more data (x-axis of Figure 3). Participants look at data to remove when it is clearly malicious data and likely hurts model learning, for instance mislabelled (#3) or outdated, historical data (#5). Additionally, we find that participants look for data that they wish to change, for example filling in or imputing missing data (#2), or correcting wrong data scales (#4) and labels (#3). Especially when data is hard to obtain (e.g. in the medical domain), removing data is not an option and correcting data is preferred (P6). Another action is the reweighting of training samples and collecting more data related to the error. These actions emphasise specific samples in the training process to encourage the model to learn them. This is particularly relevant to learning outliers (#6) and handling distribution shifts (#1).
Removal, change and addition of data have different effects on a model and therefore represent different quantifications of attribution: For instance, removal as in traditional TDA (Equation 1 in § 3) or by addition of data. Essentially, these examples map to the questions: How would the model behaviour change if one were to (a) exclude a training sample? or (b) add more training samples like the test sample to the dataset? By understanding preferences in this axis, we aim to identify what is needed for data attribution to be actionable and reflect the debugging process.
@@ 4.3.2. Axis 2: Metric
The second axis corresponds to the sample-wise metrics participants used when referring to model behaviour: the loss, the probability of the true class label, and the probability of the predicted class label (y-axis of Figure 3). The interviews show not all participants think of model behaviour in terms of sample loss, but also in classification probability. While high loss can be an indicator for wrong or outlier samples (#3, #6), participants voiced their interest in samples where the model has high or low classification probability (#7, #8). Such samples provide context of the global model behaviour to the model developer, and can therefore be helpful for their work.
The metric can be seen as a measure by which participants primarily perceive model behaviour in the debugging process. By understanding metric preferences, we aim to identify the metric that helps model developers most in debugging, testing whether the loss as in traditional TDA is the most actionable measure.

### v2

@@ 4.1.1. Participants
To get a realistic impression of user needs for TDA explanations in practice, we target participants who develop ML applications outside of academia. Our inclusion criteria were: Participants should (1) have at least one year of experience in developing ML systems and (2) work in a high-risk application area according to the EU AI Act (Parliament, 2023) (e.g., health care, cars, law enforcement. Complete list in Appendix A). The first criterion ensures that participants have hands-on experience in developing machine learning models which their statements can be rooted in. The latter criterion serves to find participants who are likely to use explanations, as these areas are subject to further regulations (Guidotti et al., 2018; Doshi-Velez and Kim, 2017). Recruiting participants poses a challenge, especially in high-risk application areas. Hence, we used purposive sampling (Guest et al., 2023) and approached participants from the authors’ network 111As the interview study is of exploratory nature, we stopped recruitment for the interviews after three months. We note that potential participants in legal and finance domains were unable to take part in the study due to non-disclosure constraints.. We recruited six participants from various domains and degrees of experience as summarised in Table 1.
| ID | Location | Domain | Job experience / with ML |
| I1 | US | Autonomous vehicles | 2 years / 7 years |
| I2 | NL | Telecommunications | 3 years / 5 years |
| I3 | FI | Computer vision for automation | 4 years / 6 years |
| I4 | NL | Health | 1 year / 3 years |
| I5 | BE | Health | 2 years / 6 years |
| I6 | PK | Health | 5 years / 2 years |
@@ 4.1.2. Interview process
The interviews were conducted in person or remotely through video call using Zoom222https://zoom.us/. All interviews were one-on-one conversations. The participants were first informed about the purpose of the study and data processing. Upon receiving informed consent, we began the interview recording. The interviews lasted between 30 and 60 minutes. Each interview addressed the following topics:
Job Roles and Responsibilities. Perspectives may vary between different domains, levels of seniority, and experience. This information provides context about the interview responses for subsequent analysis.
ML Model Development Workflow. By asking about interviewees’ development workflows, we aim to uncover diverse development styles and the challenges they face, both explicit and latent. This helps surface underlying needs that may not be readily articulated, but could inform opportunities for supporting these workflows through TDA tools.
Uses and Relevance of Training Data. We explicitly asked participants about the role that training data plays in their day-to-day tasks. Their responses highlight the types of data-centric information they find most relevant, and the specific tasks or decision points where such information becomes critical.
Perspectives on XAI and TDA. We address the participant’s perspectives on XAI and particularly on TDA to understand current explanation usage scenarios in model development and explore participant’s opinions about TDA, a data-centric approach to XAI.
@@ 4.1.3. Data analysis
The interviews were transcribed using Whisper (Radford et al., 2023) and translated using DeepL333https://www.deepl.com/translator if necessary. Then they were manually cleaned up and pseudonymised by the authors. The transcripts were analyzed through an inductive thematic analysis (Braun and Clarke, 2012) by two coders, where the codes were directly annotated in the transcript. The analysis was iterative: One interview transcript from a pilot interview was first jointly analysed in an initial coding workshop that resulted in an initial set of themes. Afterwards, the coders independently code three transcripts at a time, expanding on the themes found in the initial analysis. During an intermediate coding workshop, agreements and disagreements between the coder’s themes were discussed. The workshop resulted in a merged definition of the themes that were used for the remaining transcripts. At the intermediate coding workshop, the inter-rater agreement measured by Krippendorff’s (Krippendorff, 2011) was computed with (Marzi et al., 2024). The final coding workshop took place after both coders have reviewed the remaining transcripts. The final inter-rater agreement was .
4.2. Interview Study Findings
The analysis resulted in themes within three key meta-themes: the main challenges in working with ML systems, whether and how explanation tools are used, and perspectives and opinions about data-centric explanations like TDA. The meta-themes encompass themes and are depicted in Figure 3. We elaborate on each meta-theme below:
4.2.1. Challenges in ML development
In the interviews, we asked participants about challenges in their development workflow to explore whether XAI can address them. We identified three TDA-related themes under this meta-theme with several challenges: Data quality issues are often the root cause of model malfunction (Theme: data quality) (I1, I3, I4, I5, I6): “the models can only be as good as the […] data that you feed in.” (I3). For instance, participants encountered difficulties developing models due to missing data or absent labels (I1, I3, I4, I5, I6). Our interviews support previous observations that a significant part of model development is data work (Sambasivan et al., 2021; Jin et al., 2024). Besides data quality, our interviews revealed that the evaluation and quantification of ML systems’ performance poses a challenge in the development pipeline as well (Theme: Evaluation). In practice, developers need to balance different desiderata and objectives from their businesses and stakeholders. For example, the priority of predictive accuracy versus efficiency varies according to the particular business scenario and application. However, a large challenge in evaluation is also linked to data quality since evaluation datasets in practice are changing across time and tasks, so that missing data and unreliable data labels are a common issue. Also, the stochastic nature of ML models (I1) are pain points for the development and evaluation of ML systems since bug fixes may not work in the way they were intended, requiring further analysis (Theme: Stochasticity). Due to this stochasticity, participants perceive additional data collection as a more reliable bug fix: “[If] you have a neural network system, if you think you’ve solved the problem, you’re never 100% guaranteed […] In practice, […] whenever we had an issue […] we would first like get a lot more data, label a lot more data and add those new examples to the training data, train a new model on these.” (I1).
4.2.2. Use of (feature attribution) XAI
@@
This meta-theme groups two themes about the use or lack of XAI in practice. Our analysis revealed that, in most cases, when developers claimed to have used XAI, they were referring to feature attribution-based techniques (I1, I3, I5) (Theme: XAI-supported development). We found that feature attribution-based XAI offers explanations for per-example debugging or acts as a sanity check for model reasoning (I5, I6). Participants also use XAI to understand real-world phenomena modelled by their ML system: “The valuable part is that this model that can predict […] can also show us why […] when we use some explainability techniques.” (I2) and as a communication means with their business counterparts or to get customer buy-in for their ML systems: “[We] output [the explanation] to the business basically. That’s what they work with.” (I2), “[It’s] all about the buy-in that you get, and SHAP definitely helps.” (I5) (Theme: Reasoning checks). While explanations have different purposes, we note that participants use XAI tools mainly as an out-of-the-box functionality (e.g., the SHAP library (Lundberg and Lee, 2017)). We find that implementation thresholds must be low for the adoption of XAI in practice.
4.2.3. Perspectives on TDA
@@
None of the participants were familiar with TDA, highlighting a gap between research and practice and emphasising that TDA technology is in its early stages. Our analysis, based on discussions about the concept of TDA, identifies three themes reflecting participants’ views on XAI and TDA. Overall, participants find XAI useful for debugging and communication, though they agree its usefulness is use-case-dependent and not always guaranteed. They expect TDA not to be an exception (I2, I5) but are optimistic about its potential, especially as data is seen as the most important variable in model development (I1, I2, I3, I5): “[What] drives your model is your data. […]” (I5) (Theme: Data-focused debugging). In line with this theme, all participants also agreed that the model debugging process is largely data-focused, with training data being viewed as a lever for model performance: “It’s often the most, it’s often the biggest lever that we have in solving a problem.” (I1) (Theme: Data-focused debugging). Participants were curious whether TDA could be a potential tool for understanding bugs and enabling them to curate their datasets effectively. When comparing TDA to other XAI approaches like feature attribution (FA), participants point out that the TDA and FA target insights about model behaviour at different granularities (I2, I5, I6) and could be complimentary (Theme: Complimentary to FA). For instance, I1 mentioned that TDA could save time by identifying faulty training data while I2 pointed out that TDA and FA could be used together to identify the relationship between the faulty training data and learned features. We observed that participants have different notions of interesting training samples, with some focused on finding faulty data (I1) and others on identifying samples similar to a specific one (I5, I6) (Theme: Intuitions on data explanations).
In conclusion, our interview study highlights that TDA has significant potential to address persistent data-centric challenges in model development, particularly in handling training data quality and data-focused debugging. However, participants revealed divergent mental models of what TDA entails and what they consider as a relevant sample to identify. This suggests a conceptual gap that may hinder effective adoption of TDA-based data-centric explanations. This motivated us to further investigate which specific forms of data-centric information practitioners actually need and value during model development. In the second stage of our needfinding study, we explore these information needs to better align TDA-driven tools with real-world development workflows.
@@ 4.3. Defining a design space of TDA explanations
Drawing from our interview findings, we explore the design space of TDA explanations in model debugging. Interview participants provided rich information about what they usually look for in the training data when debugging a model, and for which purpose. Even though the objective and desired approach are similar, i.e., improving the model performance given a particular bug and data-centred debugging, respectively, developers tend to seek different information first. Specifically, we found that different data artifacts are investigated using different actions, with users interested in various metrics and considering different numbers of training samples depending on the context. From these observations, we define a three-dimensional design space (see Figure 4). In the following, we detail how each axis was driven from the themes and codes found in the interview study (Figure 3).
@@ 4.3.1. Axis 1: Action defining relevance
The first axis corresponds to the actions participants propose for different types of data artefacts: Removing data, changing the label, reweighting data and collecting more data (Figure 4). This axis is inspired by the participants’ statements about the data-centric nature of development work (Theme: Data-focused debugging). Dataset curation and manipulation is a common activity in development work and has been described as the “[largest lever]” (I1) for model performance. Different actions were mentioned or are implied in the interview study, which are linked to different hypotheses about the cause:
Participants look at data to remove when they believe that the model may have learnt the behaviour of interest from malicious data, for instance, mislabelled or outdated, historical data (Theme: Data-focused debugging Code: Artifacts are a common reason for errors, Theme: Data quality Codes: Wrong labels, Historical data). Additionally, we find that participants look for data that they wish to change, for example filling in or imputing missing data, or correcting wrong data scales and labels (Theme: Data quality Code: Data unavailability, Data scaling, Theme: Evaluation Code: Missing or unreliable data labels). Especially when data is hard to obtain (e.g. in the medical domain), removing data is not an option and correcting data is preferred (P6). Another action is the reweighting of training samples and collecting more data related to the error which participants mentioned to be rather easy in some domains like autonomous driving (P1). These actions emphasise specific samples in the training process to encourage the model to learn them (Theme: Data-focused debugging Code: New training data is easy to acquire). This is particularly relevant to learning outliers and handling distribution shifts (Theme: Data quality Code: Distribution shift, Theme: Intuitions on data explanations Code: Outlier data points).
Removal, change and addition of data have different effects on a model and therefore represent different quantifications of attribution that answer different questions. For instance, the removal action implies a question like “How would the model behaviour change if one were to exclude a training sample?” which corresponds to traditional TDA (Equation 1 in § 3). The addition of training data, however, implies the question of “How would the model behaviour change if one were to expand the dataset?” which is a largely different question. By studying preferences in this axis, we aim to identify what data attribution needs to represent to be actionable and reflect the debugging process.
@@ 4.3.2. Axis 2: Metric
The second axis corresponds to the sample-wise metrics participants used when referring to model behaviour: the loss, the probability of the true class label, and the probability of the predicted class label (y-axis of Figure 4). This axis is inspired by the quantities participants talked about when explaining their intuitions about TDA (Theme: Intuitions on data explanations).
The interviews show that not all participants think of model behaviour in terms of sample loss, but also in classification probability. While high loss can be an indicator for wrong or outlier samples (“[Samples with high training loss are] often […] mislabeled examples.” (I1), Theme: Intuitions on data explanations Code: Outlier data points), participants also talked about relevant samples as those where the model has high or low classification probability. Such samples provide context to the global model behaviour to the model developer, specifically indications about the decision boundary (“I would be interested in two kind of samples, the ones that [the model is] super confident with. But also the samples that are, I guess 50-50, it could be either one class [to see] why [the model] still chose class A.” - I4, Theme: Intuitions on data explanations Codes: Decision boundary samples, High-confidence samples) and can therefore be helpful for their work.
The metric can be seen as a measure by which participants primarily perceive per-example model performance in the debugging process: The loss is indicative of the performance of the task, but also prediction probability interpreted as model confidence in a decision matters for understanding model behaviour. Both offer different types of information about the model. By understanding metric preferences, we aim to identify the metric that helps model developers most in debugging and testing whether the loss, as in traditional TDA, is the most actionable measure.